OpenStudy (anonymous):

(cos Θ − cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@GreenCat

OpenStudy (anonymous):

@AndrewTheCookie

OpenStudy (anonymous):

@g

OpenStudy (anonymous):

@anybody

Directrix (directrix):

What are the instructions for this problem?

OpenStudy (anonymous):

To simplify it

OpenStudy (mathstudent55):

Start by squaring each binomial. Then combine like terms.

OpenStudy (mathstudent55):

To square each binomial, you can use the formulas below, or if you prefer, use FOIL. \((a + b)^2 = a^2 + 2ab + b^2\) \((a - b)^2 = a^2 - 2ab + b^2\)

OpenStudy (mathstudent55):

Wait a minute.

OpenStudy (anonymous):

Wouldn't you first add the terms inside of the parenthesis. So cos - cos = 0 and cos + cos = 2cos

OpenStudy (mathstudent55):

Are you sure it's all cosines?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

If so, then it's much simpler. On the left, cos theta - cos theta is zero. On the right cos theta + cos theta = 2 cos theta. Then square that.

OpenStudy (anonymous):

4(cos)^2 theta

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Thanks!